OpenStudy (thecuteone25):

Please help will MEDAL who ever helps the most Find the interest if the principal is $500, the interest rate is 3%,and the time is 5 years

OpenStudy (thecuteone25):

@Juliusmattos12 please help

OpenStudy (anonymous):

lemmi see.....

OpenStudy (anonymous):

I don't get it, wat is it asking

OpenStudy (thecuteone25):

I have no idea but it does saying to use the simple intrest formula I=Prt to solve

OpenStudy (thecuteone25):

@bjb2000

OpenStudy (youngstudier):

"Investment problems usually involve simple annual interest (as opposed to compounded interest), using the interest formula I = Prt, where I stands for the interest on the original investment, P stands for the amount of the original investment (called the "principal"), r is the interest rate (expressed in decimal form), and t is the time." ~ http://purplemath.com/modules/investmt.htm

OpenStudy (thecuteone25):

ok thanks

OpenStudy (youngstudier):

Do you know how to solve the problem now, @TheCuteOne25 ? ^-^

OpenStudy (thecuteone25):

yes that helped a lot thanks @YoungStudier

OpenStudy (youngstudier):

No problem. ~(^_^)~

OpenStudy (wolf1728):

I always thought these problems almost always involved compound interest. The formula is: Total = Principal * (1 + rate)^ years Total = 500.00 * (1.03) ^ 5 Total = 500.00 * 1.1592740743 Total = 579.64 So interest is $79.64

OpenStudy (thecuteone25):

oh ok thank you

OpenStudy (wolf1728):

Oh, and here's a link to the information and formula I used: http://www.1728.org/compint.htm

OpenStudy (thecuteone25):

thanks

OpenStudy (wolf1728):

u r welcome :-)